CWUIN. IY 1 Score oft Section 14.1 Exercise 1 HW Score: SON, 2 of 4 pts Question
Answer #1

the necessary calculation table is :-

birth time observed expected \frac{(obs-exp)^2}{exp}
winter 25 120/4 = 30 (25-30)2/30 = 0.83333
spring 37 120/4 = 30 1.63333
summer 31 120/4 = 30 0.03333
fall 27 120/4 = 30 0.3
total 120 2.79999

a).expected number of birth in each season is = 30

b).the chi square statistic is ;-

\chi^2=\sum\frac{(obs-exp)^2}{exp}=2.79999 = \mathbf{2.80}

c).the degrees of freedom is = (4-1 )= 3

***in case of doubt, comment below. And if u liked the solution, please like.